OpenStudy (anonymous):

boat is pushed to water. it moves distance s in time t with formula: s = 25m*(1-e^-(k*t)) where k is boat constant. when time is 6s boat has moved 10m. what is distance after 26 seconds? I don't know how to remove exponent. can someone help me to right direction? :/ 10m = 25*(1 - e^-(k*6)) //:25 0.4 = 1 - e^-(k*6) // -1 -0.6 = e^-(k*6) ln(k*6) = ln(0.6) so something is wrong.

OpenStudy (agent0smith):

Your last step isn't right. You also dropped a negative. \[-0.6 = -e ^{-(6k)}\] Multiply it all by -1, then taking logs of both sides yields \[\ln 0.6 = {-(6k)}\] This is the log property you messed up on: \[\ln e^x = x ln e = x\] but you had done it as \[\ln e^x = \ln x\]

OpenStudy (anonymous):

so ln(0.6)\[k = \frac{ \ln(0.6) }{ 6 } = -0.085137603\] i think that's right. thanks :)

OpenStudy (agent0smith):

Almost, but you dropped a negative again. \[\ln 0.6 = {-(6k)} \] \[k = \frac{ - \ln 0.6 }{ 6 }\] So your answer should be +0.0851